### v1

@@ 1 Introduction
Aortic aneurysm, which is morphologically defined as the focal dilation and structural degradation of the aorta, is an asymptomatic disease, and its rupture is a significant cause of death worldwide. For instance, in the United States of America, aortic aneurysms and dissections cause over $10,000$ deaths each year (Go et al., 2013; Milewicz and Ramirez, 2019). To date, thoracic endovascular aortic repair (TEVAR) is the only effective treatment for aortic aneurysm. TEVAR has been a particularly attractive surgical intervention because it is less invasive than conventional open surgical repair (Goto et al., 2019; Narita et al., 2016; Shijo et al., 2016; Shimamura et al., 2008).
@@
The aorta is the largest conduit artery in the body, and due to its extraordinary ability to expand and contract, it serves as a reservoir that transforms the highly pressured and pulsatile heart output into a flow with moderate fluctuations (Ku, 1997). Deterioration of this buffering effect due to stiffening of the arterial wall is the main cause of hypertension (Ku, 1997; Liu and Fung, 1989). These dynamics of the aorta are also important for propagation of the pulse wave, which has been recognized as an early indicator of the health status of the cardiovascular system (Hametner et al., 2021; Laurent et al., 2006; Townsend et al., 2015). Thus, understanding the hemodynamic differences between pre- and postoperative TEVAR patients based on pulse-wave dynamics is of paramount importance, not only for surgical decision-making to achieve optimal clinical outcomes, but also for evaluating postoperative hemodynamics.
@@
In TEVAR, the aortic arch is mapped according to the segmentation of the vertebral column or landing zone (Ishimaru, 2004), where zone 00 ranges from the ascending aorta (AA) to the origin of the innominate artery (IA), zone $1$ ranges from the IA to the origin of the left common carotid artery (LCCA), zone $2$ ranges from the LCCA to the left subclavian artery (LSA), and zones 3 and 4 follow in the longitudinal direction. TEVAR for partial arch debranching, known as debranching TEVAR (dTEVAR), includes both stent graft implantation and extra-anatomical bypass, where the stent graft is placed from zone $1$ or $2$ with a length of several millimeters depending on the patient. This operation is most often performed in high-risk patients with thoracic aortic aneurysms (TAA) (Narita et al., 2016; Shijo et al., 2016; Shimamura et al., 2008). Currently, expanded polytetrafluoroethylene (ePTFE) and polyethylene therephthalate (PET) are the two most popular graft materials in abdominal aortic aneurysm repair due to their remarkable biocompatibility and durability. Additional supra-arch vessel reconstruction in dTEVAR is known to prevent cerebral infarction (Hajibandeh et al., 2016; Shigemura et al., 2000; Ullery et al., 2012; Yoshitake et al., 2016) and ischemia in the cerebral circulation (Cooper et al., 2009; Feezor and Lee, 2009; Patterson et al., 2014).
@@
A previous experimental study has shown that total intracranial blood flow was preserved after dTEVAR, with significant decrease of flow in the LVA and significant increase of flow in the RVA (Goto et al., 2019). The hemodynamic mechanism of such changes in posterior cerebral circulation after dTEVAR remains uncertain yet due the technical difficulty of measurements. Thus, our primary concern here is to evaluate bilateral vertebral artery blood flows under pulsation between the patients before and after $1$dTVEAR with axillo-axillary artery (AxA-AxA) bypass, and to clarify the hemodynamic mechanism associated with postoperative significant changes of posterior cerebral circulation. Since dTEVAR includes extra-anatomical bypass, it was expected that the intracranial hemodynamic changes due to dTEVAR would be coupled with fluid flow and pulse waves in the VA. Therefore, the objective of this study was to numerically clarify whether there exist phase differences of the flow rate in the VA, as a representative intracranial vessel, between patients before and after dTEVAR, especially in cases with one bypass from the right axillary artery (RAxA) to the left axillary artery (LAxA) and embolization at the branch point between a stent-grafted aortic arch and the LSA (see Figure 1). If phase differences of the flow rate exist, we clarify how surgery contributes to it. Hereafter, we refer to this operation as single-debranching TEVAR ($1$dTEVAR). A one-dimensional ($1$D) model analysis was used to numerically investigate blood flow rates in the L/RVA in pre- and postoperative patients after $1$dTEVAR.
@@
In a human aorta with $\sim 10$ mm radius and $\sim 0.3$ mm thickness, the wave speed, quantified by the traditional foot-to-foot wave velocity method, e.g., by Laurent et al. (2006), is over $5$ m/s (Latham et al., 1985). Such conventional experimental methodology, however, is still difficult to quantify the effect of pulse waves in the VA on the intracranial blood supply. Furthermore, assuming that one pulse has a duration of $1$ second, the ratio between the wave length and radius is $500$, i.e., long wave. Numerical analysis for blood flow with long-wave pulses still involves a heavy computational load even with a two-dimensional model, although several attempts for this issue were reported, e.g., in Yokoyama et al. (2021). $1$D model analysis is one of the most effective and practical non-invasive solutions to investigate hemodynamics with long-wave pulses (Formaggia et al., 2003; Taylor and Draney, 2004; van de Vosse and Stergiopulos, 2011). Thus, for practical purposes, 1D modeling has been widely used in circulatory systems, such as the systemic arteries (Alastruey et al., 2011; Olufsen, 1999), coronary circulation (Duanmu et al., 2019; Huo and Kassab, 2007; Mynard et al., 2014), the circle of Willis (Alastruey et al., 2007), and large vasculature, including venous systems coupled with contributions of the microcirculation modeled by lumped parameters (Alastruey et al., 2008; Duanmu et al., 2019; Huo and Kassab, 2007; Liang et al., 2009; Mynard et al., 2012, 2014). Recently, $1$D modeling was applied to a problem involving both blood flow rate and blood oxygenation (Feiger et al., 2020). Padmos et al. (2020) used patient-specific medical imaging data on the anatomy of the circle of Willis to perform a $1$D model analysis of intracranial blood flow assuming a steady state, coupled with three-dimensional tissue perfusion. In this study, the experimental evidence on differences in the flow rate in the L/RVA between patients before and after $1$dTEVAR was reported. Next, using 1D numerical model, the phase difference of the flow rate in the L/RVA in pre- and postoperative patients was evaluated. Simulations were also performed for different local vessel stiffnesses and lengths of vessels that had stiffened due to stent grafting.
@@ 2.1 Subjects and measurements
This retrospective study was conducted in accordance with the guidelines of the Declaration of Helsinki. All experimental protocols were approved by the institutional review board of Osaka University. All subjects provided oral and written informed consent to participate in this study. A prospective analysis of blood flow in the left and right internal carotid arteries (L/RICA) and the L/RVA in 9 patients before and after dTEVAR was performed between January $2015$ and January $2020$ at Osaka University Hospital, where RAxA-LAxA bypass was performed (see also Figure 1B). A ringed $8$-mm ePTFE graft (FUSION, MAQUET Getinge Group, Japan) was used in all the debranching procedures (see also (Goto et al., 2019)). Two-dimensional ($2$D) cine phase-contrast data was acquired with $3$T magnetic resonance imaging (MRI) (MR750750750750-$3$T, GE Healthcare, Waukesha, WI, USA). Validations of the resolution of $3$T MRI have been performed in in vivo studies, for instance, by Calderon-Arnulphi et al. (2011) and Zhao et al. (2007). In all patients, MRI was performed within $6$ months prior to the dTEVAR procedure and within $1$ month afterward (Goto et al., 2019).
Table 1: Geometrical parameters ($r_{0}$ and $L$) in a representative patient
. ID Artery Radius $r_{0}$ [mm] Length $L$ [mm] 00 Left Ventricle (LV; inlet boundary) $12.5$ $1$ Ascending Aorta (AA) $12.5$ $50$ $2$ Aortic Arch I $12.5$ $20$ $3$ Aortic Arch II $12.5$ $10$ $4$ Thoracic Aorta $12.5$ $30$ $5$ Brachiocephalic $7$ $33$ $6$ Right Subclavian (RSA) $4$ $54$ $7$ Right Axillary $4$ $139$ $8$ Right Vertebral (RVA) $1.5$ $200$ $9$ Right Carotid $2.5$ $100$ $10$ Right External Carotid (RECA) $2$ $50$ $11$ Right Internal Carotid (RICA) $2$ $100$ $12$ Left Carotid $2.5$ $130$ $13$ Left External Carotid (LECA) $2$ $50$ $14$ Left Internal Carotid (LICA) $2$ $100$ $15$ Left Subclavian (LSA) $4$ $60$ $16$ Left Axillary $4$ $150$ $17$ Left Vertebral (LVA) $1.5$ $200$- Bypass $4.5$ $262$- Stent-Grafted Region $12.5$ $40$ (–$240$)
@@
The vascular tree models of both pre- and postoperative patients were reconstructed based on geometrical parameters of vessel diameter and length between branch points, as shown in Figure 1. Arterial vascular geometries, including reference radius ($r_{0}$) and length ($L$) in a representative patient, were measured both pre- and postoperatively at Osaka University Hospital, and the data are summarized in Table 1. The diameters of the end terminals, which are denoted by the vessel ID $=\chi^{\prime}$ in Figure 1(A), were determined so that the time-average flow rate became similar to that obtained with experimental measurements by Bogren and Buonocore (1994) or Goto et al. (2019) (see Figure 4). The lengths of the end terminals were uniformly set to $50$ mm. For simplicity, branch angles were uniformly set to $30$ deg. The effect of branch angle is mentioned in § 3.5. To represent arterial trees in patients after $1$dTEVAR, The Young’s modulus was set to be almost $100$ times larger in the stent-grafted area (vessel ID $=4$; see Table 1) than in preoperative patients, and also created a bypass from the RAxA to the LAxA (vessel ID $=7$ and $16$) with embolization at the branch point from the stent-grafted thoracic aorta to the LSA (see Figure 1(B)). In this study, the bypass was considered to have a radius of $4.5$ mm and a length of $262$ mm, and the stent-grafted artery was considered to have a radius of $12.5$ mm and a length of $40$ mm. The diameters and lengths of end terminals were also the same as those in the preoperative vascular tree model.
@@ 2.2 Mathematical model and simulation
The flow and pulse wave are integrated over time by the Lax-Wendroff method with $\Delta t=10^{-5}$ s from $t^{n}$ to $t^{n+1}=t^{n}+\Delta t$. The precise descriptions are given in the Appendix.
@@
$\alpha$ Coefficient of the velocity profile $1$ Sherwin et al. (2003)
@@
To capture the internal pressure profile in the heart during a cardiac cycle, wherein vascular pumping is accelerated in the systolic phase and attenuated in the diastolic phase (Morishita et al., 2021), the inlet pressure $p_{\mathrm{in}}(t)$ is given as (Yokoyama et al., 2021)
@@
where $p_{a}$ ($=4333$ Pa) is the amplitude from the base pressure $p_{0}$ ($=10666$ Pa), $T$ ($=1$ s) is the wave period, and $\epsilon_{\mathrm{s}}$ or $\epsilon_{\mathrm{d}}$ is the waveform parameter (black line in Figure 2). Hereafter, the subscripts s and d denote the systolic and diastolic phases, respectively. In this study, $\epsilon_{\mathrm{s}}=0.1$ and $\epsilon_{\mathrm{d}}=0$ (Yokoyama et al., 2021) were used, so that the derivatives up to the fourth of $p(t)$ are continuous. The waveform reflects a fast expansion during the systolic phase $40\leq t<T/$ 4 and a slow contraction during the diastolic phase $T/4\leq t<T$. The parameters in our simulations and their values are shown in Table 2.
Figure 2 shows pressures and calculated flow rates during a period $T$ ($=1$ s). The model pressure given in equations equation (7) and (8) well captures the profile of human aortic pressure (Figure $3$a in the work by van de Vosse and Stergiopulos (2011)). A no-reflection condition was applied for each outlet by a backward Riemann invariant $W_{-}=0$. A more precise description of the methodology is presented in the Appendix.
@@ 3.2 Model validation
Using the same preoperative vascular tree model, the flow rate ratio between the L/RICA and L/RVA was calculated. The flow rate ratio in patients after $1$dTEVAR was also calculated with the postoperative vascular tree model (Figure 1B). The diameters of the end terminals were the same as those in the preoperative vascular tree model. The calculated flow rate ratio in each artery was compared with previous experimental measurements by Goto et al. (2019) (Figures 4B and 4C). In the experiments, although there were no large differences before and after surgery in the flow rate ratio in the L/RICA or the LVA, the flow rate ratio in the RVA increased after surgery. The numerical results qualitatively agree with these experimental measurements.
@@ 3.3 $1$D model analysis of the flow rate in the ICA in pre-and postoperative patients
To gain insight into the mechanism of the phase difference in the LVA and the lack of such a difference in the RVA (Figure 3), $1$D numerical simulations were performed using arterial vascular tree models before and after surgery (Figure 1). Model verifications are shown in the Appendix (see Figures A2). The time history of the pressure and flow rates in the L/RVA during a cardiac cycle $T$ are shown in Figure 5(A) and 5(B), respectively, where the results before and after surgery are superposed. Data are shown after the pressure and flow rates have reached the stable periodic phase ($t\geq 2$ s). As described regarding the experimental measurements of flow rates (Figure 3(A)), the numerical results of the flow rate waveform in the LVA were delayed, by $6.8$% of a single period, which is similar to the experimental measurements (Figure 5(A)). As in experimental measurements, the phase difference was also quantified by the time point observed at the maximum flow rate. On the other hand, the numerical results of the flow rate waveform in the RVA were only slightly delayed, by $0.84$% of a single period (Figure 5(B)). We concluded that our numerical results of the phase difference in the L/RVA are consistent with the experimental measurements (Figure 3(B)).

### v2

@@ 1 Introduction
Aortic aneurysm, which is morphologically defined as the focal dilation and structural degradation of the aorta, is an asymptomatic disease, and its rupture is a significant cause of death worldwide. For instance, in the United States of America, aortic aneurysms and dissections cause over $10,000$ deaths each year [1, 2]. To date, thoracic endovascular aortic repair (TEVAR) is the only effective treatment for aortic aneurysm. TEVAR has been a particularly attractive surgical intervention because it is less invasive than conventional open surgical repair [3, 4, 5, 6].
@@
The aorta is the largest conduit artery in the body, and due to its extraordinary ability to expand and contract, it serves as a reservoir that transforms the highly pressured and pulsatile heart output into a flow with moderate fluctuations [7]. Deterioration of this buffering effect due to stiffening of the arterial wall is the main cause of hypertension [7, 8]. These dynamics of the aorta are also important for propagation of the pulse wave, which has been recognized as an early indicator of the health status of the cardiovascular system [9, 10, 11]. Thus, understanding the hemodynamic differences between pre- and postoperative TEVAR patients based on pulse-wave dynamics is of paramount importance, not only for surgical decision-making to achieve optimal clinical outcomes, but also for evaluating postoperative hemodynamics.
@@
In TEVAR, the aortic arch is mapped according to the segmentation of the vertebral column or landing zone [12], where zone 00 ranges from the ascending aorta (AA) to the origin of the innominate artery (IA), zone $1$ ranges from the IA to the origin of the left common carotid artery (LCCA), zone $2$ ranges from the LCCA to the left subclavian artery (LSA), and zones 3 and 4 follow in the longitudinal direction. TEVAR for partial arch debranching, known as debranching TEVAR (dTEVAR), includes both stent graft implantation and extra-anatomical bypass, where the stent graft is placed from zone $1$ or $2$ with a length of several millimeters depending on the patient. This operation is most often performed in high-risk patients with thoracic aortic aneurysms (TAA) [4, 5, 6]. Currently, expanded polytetrafluoroethylene (ePTFE) and polyethylene therephthalate (PET) are the two most popular graft materials in abdominal aortic aneurysm repair due to their remarkable biocompatibility and durability. Additional supra-arch vessel reconstruction in dTEVAR is known to prevent cerebral infarction [13, 14, 15, 16] and ischemia in the cerebral circulation [17, 18, 19].
@@
A previous experimental study has shown that total intracranial blood flow was preserved after dTEVAR, with significant decrease of flow in the LVA and significant increase of flow in the RVA [3]. The hemodynamic mechanism of such changes in posterior cerebral circulation after dTEVAR remains uncertain yet due the technical difficulty of measurements. Thus, our primary concern here is to evaluate bilateral vertebral artery blood flows under pulsation between the patients before and after dTVEAR with axillo-axillary artery (AxA-AxA) bypass, and to clarify the hemodynamic mechanism associated with postoperative significant changes of posterior cerebral circulation. Since dTEVAR includes extra-anatomical bypass, it was expected that the intracranial hemodynamic changes due to dTEVAR would be coupled with fluid flow and pulse waves in the VA. Therefore, the objective of this study was to numerically clarify whether there exist phase differences of the flow rate in the VA, as a representative intracranial vessel, between patients before and after dTEVAR, especially in cases with one bypass from the right axillary artery (RAxA) to the left axillary artery (LAxA) and embolization at the branch point between a stent-grafted aortic arch and the LSA (see Figure 1). If phase differences of the flow rate exist, we clarify how surgery contributes to it. Hereafter, we refer to this operation as single-debranching TEVAR ($1$dTEVAR). A one-dimensional ($1$D) model analysis was used to numerically investigate blood flow rates in the L/RVA in pre- and postoperative patients after $1$dTEVAR.
@@
In a human aorta with $\sim 10$ mm radius and $\sim 0.3$ mm thickness, the wave speed, quantified by the traditional foot-to-foot wave velocity method, e.g., by Laurent et al. [10], is over $5$ m/s [20]. Such conventional experimental methodology, however, is still difficult to quantify the effect of pulse waves in the VA on the intracranial blood supply. Furthermore, assuming that one pulse has a duration of $1$ second, the ratio between the wave length and radius is $500$, i.e., long wave. Numerical analysis for blood flow with long-wave pulses still involves a heavy computational load even with a two-dimensional model, although several attempts for this issue were reported, e.g., in [21]. $1$D model analysis is one of the most effective and practical non-invasive solutions to investigate hemodynamics with long-wave pulses [22, 23, 24]. Thus, for practical purposes, 1D modeling has been widely used in circulatory systems, such as the systemic arteries [25, 26], coronary circulation [27, 28, 29], the circle of Willis (CoW) [30], and large vasculature, including venous systems coupled with contributions of the microcirculation modeled by lumped parameters [27, 28, 29, 31, 32, 33]. Recently, $1$D modeling was applied to a problem involving both blood flow rate and blood oxygenation [34]. Padmos et al. [35] used patient-specific medical imaging data on the anatomy of the CoW to perform a $1$D model analysis of intracranial blood flow assuming a steady state, coupled with three-dimensional tissue perfusion. In this study, the experimental evidence on phase differences in the flow rate in the L/RVA between patients before and after $1$dTEVAR was reported. Next, using 1D numerical model, the phase difference of the flow rate in the L/RVA in pre- and postoperative patients was evaluated. Simulations were also performed for different local vessel stiffnesses and lengths of vessels that had stiffened due to stent grafting.
@@ 2.1 Subjects and measurements
This retrospective study was conducted in accordance with the guidelines of the Declaration of Helsinki. All experimental protocols were approved by the institutional review board of Osaka University. All subjects provided oral and written informed consent to participate in this study. A prospective analysis of blood flow in the left and right internal carotid arteries (L/RICA) and the L/RVA in 9 patients before and after dTEVAR was performed between January $2015$ and January $2020$ at Osaka University Hospital, where RAxA-LAxA bypass was performed (see also Figure 1B). All patients were male, with a mean age of $73.7\pm 14.4$ years (standard deviation, SD). The characteristics of the enrolled patients are presented in Table 1. Note that a magnetic resonance imaging (MRI) confirmed no stenosis in the CoW in preoperative patients. A ringed $8$-mm ePTFE graft (FUSION, MAQUET Getinge Group, Japan) was used in all the debranching procedures (see also [3]). Two-dimensional ($2$D) cine phase-contrast data was acquired with $3$T MRI (MR750750750750-$3$T, GE Healthcare, Waukesha, WI, USA). Validations of the resolution of $3$T MRI have been performed in in vivo studies, for instance, by Calderon-Arnulphi et al. [36] and Zhao et al. [37]. In all patients, MRI was performed within $6$ months prior to the 1dTEVAR procedure and within $1$ month afterward [3].
Table 1: Patient characteristics.
Number of patients, $n$ $9$
Mean age (SD) $73.7(14.4)$
Gender (men::::women) $9:0$
[…]
         Cardiovascular disease $6/9(66.7\%)$
         COPD $2/9(22.2\%)$
         Chronic kidney disease $3/9(33.3\%)$
         Haemodialysis $0/9(0\%)$
         Diabetes mellitus $2/9(22.2\%)$
         Hypertension $9/9(100\%)$
         Dyslipidaemia $5/9(55.6\%)$
         Cerebrovascular disease (Old cerebral infarction) $2/9(22.2\%)$
COPD: Chronic obstructive pulmonary disease
The vascular tree models of both pre- and postoperative patients were reconstructed based on geometrical parameters of vessel diameter and length between branch points, as shown in Figure 1. Arterial vascular geometries, including reference radius ($r_{0}$) and length ($L$) in a representative patient, were measured both pre- and postoperatively at Osaka University Hospital, and the data are summarized in Table 2. The diameters of the end terminals, which are denoted by the vessel ID $=\chi^{\prime}$ in Figure 1(A), were determined so that the time-average flow rate became similar to that obtained with experimental measurements by Bogren & Buonocore [38] or Goto et al. [3] (see Figure 4). The lengths of the end terminals were uniformly set to $50$ mm. For simplicity, branch angles were uniformly set to $30$ deg. The effect of branch angle is mentioned in Appendix §A.1 (see Figure A1). To represent arterial trees in patients after $1$dTEVAR, the Young’s modulus was set to be almost $100$ times larger in the stent-grafted area (vessel ID $=4$; see Table 2) than in preoperative patients, and also created a bypass from the RAxA to the LAxA (vessel ID $=7$ and $16$) with embolization at the branch point from the stent-grafted thoracic aorta to the LSA (see Figure 1(B)). In this study, the bypass was considered to have a radius of $4.5$ mm and a length of $262$ mm, and the stent-grafted artery was considered to have a radius of $12.5$ mm and a length of $40$ mm. The diameters and lengths of end terminals were also the same as those in the preoperative vascular tree model.
Table 2: Geometrical parameters ($r_{0}$ and $L$) in a representative patient.
ID Artery Radius $r_{0}$ [mm] Length $L$ [mm]
00 Left Ventricle (LV; inlet boundary) $12.5$
$1$ Ascending Aorta (AA) $12.5$ $50$
$2$ Aortic Arch I $12.5$ $20$
$3$ Aortic Arch II $12.5$ $10$
$4$ Thoracic Aorta $12.5$ $30$
$5$ Brachiocephalic $7$ $33$
$6$ Right Subclavian (RSA) $4$ $54$
$7$ Right Axillary $4$ $139$
$8$ Right Vertebral (RVA) $1.5$ $200$
$9$ Right Carotid $2.5$ $100$
$10$ Right External Carotid (RECA) $2$ $50$
$11$ Right Internal Carotid (RICA) $2$ $100$
$12$ Left Carotid $2.5$ $130$
$13$ Left External Carotid (LECA) $2$ $50$
$14$ Left Internal Carotid (LICA) $2$ $100$
$15$ Left Subclavian (LSA) $4$ $60$
$16$ Left Axillary $4$ $150$
$17$ Left Vertebral (LVA) $1.5$ $200$
-- Bypass $4.5$ $262$
-- Stent-Grafted Region $12.5$ $40$ (–$240$)
@@ 2.2 Mathematical model and simulation
The flow and pulse wave are integrated over time by the Lax-Wendroff method with $\Delta t=10^{-5}$ s from $t^{n}$ to $t^{n+1}=t^{n}+\Delta t$. The precise descriptions are given in the Appendix §A.2.
@@
$\alpha$ Coefficient of the velocity profile $1$ [39]
@@
To capture the internal pressure profile in the heart during a cardiac cycle, wherein vascular pumping is accelerated in the systolic phase and attenuated in the diastolic phase [43], the inlet pressure $p_{\mathrm{in}}(t)$ is given as [21]
@@
where $p_{a}$ ($=4333$ Pa) is the amplitude from the base pressure $p_{0}$ ($=10666$ Pa), $T$ ($=1$ s) is the wave period, and $\epsilon_{\mathrm{s}}$ or $\epsilon_{\mathrm{d}}$ is the waveform parameter (black line in Figure 2). Hereafter, the subscripts s and d denote the systolic and diastolic phases, respectively. In this study, $\epsilon_{\mathrm{s}}=0.1$ and $\epsilon_{\mathrm{d}}=0$ [21] were used, so that the derivatives up to the fourth of $p(t)$ are continuous. The waveform reflects a fast expansion during the systolic phase $40\leq t<T/$ 4 and a slow contraction during the diastolic phase $T/4\leq t<T$. The parameters in our simulations and their values are shown in Table 3.
Figure 2 shows pressures and calculated flow rates during a period $T$ ($=1$ s). The model pressure given in equations equation (7) and (8) well captures the profile of human aortic pressure (Figure $3$a in the work by [24]). A no-reflection condition was applied for each outlet by a backward Riemann invariant $W_{-}=0$. A more precise description of the methodology is presented in the Appendix §A.2.
@@ 3.2 Model validation
Using the same preoperative vascular tree model, the flow rate ratio between the L/RICA and L/RVA was calculated. The flow rate ratio in patients after $1$dTEVAR was also calculated with the postoperative vascular tree model (Figure 1B). The diameters of the end terminals were the same as those in the preoperative vascular tree model. The calculated flow rate ratio in each artery was compared with previous experimental measurements by Goto et al. [3] (Figures 4B and 4C). In the experiments, although there were no large differences before and after surgery in the flow rate ratio in the L/RICA or the LVA, the flow rate ratio in the RVA increased after surgery. The numerical results qualitatively agree with these experimental measurements.
@@ 3.3 $1$D model analysis of the flow rate in the ICA in pre-and postoperative patients
To gain insight into the mechanism of the phase difference in the LVA and the lack of such a difference in the RVA (Figure 3), $1$D numerical simulations were performed using arterial vascular tree models before and after surgery (Figure 1). Model verifications are shown in the Appendix §A.3 (see Figures A3). The time history of the pressure and flow rates in the L/RVA during a cardiac cycle $T$ are shown in Figure 5(A) and 5(B), respectively, where the results before and after surgery are superposed. Data are shown after the pressure and flow rates have reached the stable periodic phase ($t\geq 2$ s). As described regarding the experimental measurements of flow rates (Figure 3(A)), the numerical results of the flow rate waveform in the LVA were delayed, by $6.8$% of a single period, which is similar to the experimental measurements (Figure 5(A)). As in experimental measurements, the phase difference was also quantified by the time point observed at the maximum flow rate. On the other hand, the numerical results of the flow rate waveform in the RVA were only slightly delayed, by $0.84$% of a single period (Figure 5(B)). We concluded that our numerical results of the phase difference in the L/RVA are consistent with the experimental measurements (Figure 3(B)).